1. A hypothetical closed economy has a national income model of the form y = C + I + G where C = 30 + 0.8Y and I and G and private investment and government expenditure are exogenously determined t 50 and 80 units respectively. Compute the national equilibrium level of income for this economy using aggregate income equals aggregate expenditure and withdrawals and equal injection methods.
1
Expert's answer
2020-07-29T11:32:35-0400

Y = C + I +G

                       Y = 30 + 0.8Y +50 + 80

                       Y – 0.8Y = 160

                          Y* = 160/0.2

                         Y* = 800